OpenStudy (anonymous):

how to integrate sin^2(x)cos^4(x)

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin ^{2}x \cos ^{4}x\]\[\int\limits_{}^{}(2\sin ^{2}xcos ^{4}x)/2\]\[\int\limits\limits_{}^{}(1-\cos x/2)\cos ^{4}x/4\] In the same way dissociate cos^4 x into 1-cos^2 x/2 and then into 1-cos^2 x/4 and continue its way too much to write the equations here

OpenStudy (anonymous):

1+cos^4x

OpenStudy (anonymous):

Do you need the answer or the working

OpenStudy (anonymous):

I need the working

OpenStudy (anonymous):

damn the answer the huge

OpenStudy (anonymous):

lol can you explain more, i still don't get how it works

OpenStudy (anonymous):

\[2\sin ^{2}x=1-cosx/2\] \[2\cos ^{2}x=1+cosx/2\]

OpenStudy (anonymous):

using theese equations first remove the powers

OpenStudy (dumbcow):

you could also substitute sin^2 = 1-cos^2 --> integral cos^4 - cos^6 then you have to look up the reduction formula for cos^n